ADNAN MUHAMMED

Grade 12,

Solve the following equations: 6z + 10 = - 2.........

 Solve the following equations:
 6z + 10 = - 2.........

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

By transposing 10 from LHS to RHS it becomes - 10
Then,
= 6z = -2–10
= 6z = - 12
Now,
Divide both side by 6,
= 6z/6 = -12/6
= m = -2

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
6z + 10 = -2
6z = -2-10
6z = -12
Z = -2
 
Hope it helps!
Thank you ✌️